Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Wednesday, December 23, 1998

The following is a brief summary of the environmental monitoring activities that were completed on December 23, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

12/23/98 (Wednesday)

Segment 2B (Castaic to Sylmar): Report for 12-23-98. Clean up continued along the right of way and at several extra work sites. At the Yard in Castaic, crews removed the fuel tanks and began excavating contaminated soil from the area. In addition, removal of equipment and debris progressed along the right of way. At Balboa Blvd de-watering of segment 2 continued while crews planned how to complete the electrical tie-in to the Balboa vault.

Segment 3C (Sylmar to San Fernando): Report to 12-23-98. Crews began loading the 20'' pipe with water for the hydro-test, meanwhile clean up of the right of way continued and the forms were removed from the concrete barriers next to the high speed channel. At the Tujunga wash crews compacted soil next to the bridge and clean up continued at Tuxford St. Segment 3A (San Fernando to Taylor Yard): Fiber optic crews spliced cable at Van Nuys Blvd and Buena Vista St while electrical crews worked on the vault site at Victory St.

Inspection Summar

In the past 24 hours, 2 inspection event(s) were reported for 2 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page